# Evaluating Time-Dependent Methods and Seasonal Effects in Code Technical Debt Prediction

### v1

@@ 3.6.2 Execution process for ARIMA+LM and SARIMA + LM models
3.7 Replicability
To allow verifiability and replicability, we made the raw data and analysis files available in our online appendix.666 https://figshare.com/s/61eeca3d832408c9424f
4 Results
In this section, we report the results obtained by answering our RQs. The dataset included original commit data for 31 OS Java projects and the serialization and posterior linear interpolation created data serialized in biweekly and monthly time series. We provide results for 14 software projects which were successfully processed through all the adopted models in this study. We encountered linear algebra convergence issues while fitting the TSA models. We report this issue in the model fitting stage as a threat to the validity of this study, which we formally state in Section 6 and provide further details on it later in this section. Therefore, since we encountered this issue in the execution of different models across the collected projects, and for the sake of fairness in the models’ results, we excluded projects that were not able to provide results for all the considered models in the study. To answer our RQs, we present the results of the performed analysis as the average score of the outputs obtained with each prediction model for the defined prediction performance metrics. Likewise, the aggregation step is performed for the results obtained with biweekly and monthly data separately. Figure 6 provides an illustrated example of the MAE results obtained for the executed models in each of the analyzed projects. We provide all the result figures and tables before aggregation in the shared online appendix (see Section 3.7).
4.1 Descriptive Analysis
We now present the descriptive statistics of the initially considered subjects before building our prediction models. Table 2 shows the statistical characteristics of the projects within the raw data set as well as the generated time series data. From the presented table, we discovered two opposite trends in the frequency of the collected projects used to perform SQ analyses in their code. On one side, we observed projects whose number of analyzed commits in the collected raw dataset decreased during the time serialization to achieve the monthly and biweekly datasets (e.g. felix in Table 2). This phenomenon was given in projects whose SQ analysis activity was intensive during the period of time in which the data collection was performed. On the other side, some projects presented the inverse situation as their number of observations increased during the time serialization process (e.g. fileupload in Table 2). This reflected a scarce use of of SQ during the time frame of collected observations and therefore required an intensive data interpolation action to fill the missing time observations.
At this stage of the analysis, we realized that depending on the frequency of the execution of SQ analyses, the quality of the data would vary. The presented variations suggested that performing static analysis with tools like SQ in a fixed frequency could result in concise, authentic, and well-monitored time series data on the state of the quality of the code base of projects. In contrast, less frequent or randomly executed SQ analyses provide insights into the code base throughout time but leave notable periods of time with missing data, thus requiring the action of data interpolation to perform TSA. Moreover, due to the missing data for some of the projects, the algorithms from the adopted TSA models failed to converge during the model fitting stage described in Section 3.5.1. Due to the existing complexity in the data, we encountered several linear algebra decomposition issues [55] involved in the parameter estimation process while fitting the TSA models. As mentioned earlier, we leveraged preprocessing the set of independent variables within the model training through the backward variable selection criteria as well as standardization. However, Table 3 reports the cases in which the adopted models reported linear algebra decomposition issues, which impeded obtaining results for all the models in each of the projects that reported issues respectively.
@@ 4.2 On the performance of Code TD prediction with existing time-dependent approaches (RQ1)
We conducted the prediction analysis described in Section 3.6 for each of the 31 projects in our dataset. Figure 5 shows the unified graphical illustration of the predictive performance results from the adopted performance metrics. The displayed results demonstrate a clear superiority of ARIMAX compared to ARIMA+LM with both biweekly and monthly serialized time series data. Considering the results for MAE and RMSE, the differences in prediction error between the ARIMAX and ARIMA+LM are 14,000 and 14,666 within the biweekly data, and 15,237 and 15,931 within the monthly data, thus denoting a remarkable difference among both approaches. Hence, we can affirm that given the existing TSA models for Code TD prediction, ARIMAX model overcomes the predictive performance of ARIMA+LM model. To be concise, both models performed better when being trained with biweekly data as depicted by MAE and RMSE, whereas MAPE results provided an improvement in the predictive performance for ARIMA+LM model when training it with monthly data.
We compared the ARIMAX and ARIMA+LM models against the adopted ML algorithms commonly used in the field. Tables 4 and 5 present the prediction performance results for the biweekly and monthly data, respectively. More specifically, the tables provide the average results calculated from the prediction results obtained for the 14 projects. Results for the individual projects are included in the replication package. It is worth noting that in both cases, using biweekly and monthly data, ARIMAX model demonstrated a clear superiority over the rest of the models, thus providing a more accurate prediction outcome for future Code TD observations. Results such as 1,525.96 for MAE, as well as 2,646.20 for RMSE within the biweekly results, compared to the ones presented by the next model following the ranking of results, RF, denote a quantified improvement of 717 and 1,232 in the described prediction errors accordingly. Moreover, RF proved to be the second-best prediction model. As described previously, RF assumes non-linear relationships between variables. Consequently, despite the indications that TSA models such as ARIMAX may be appropriate for predicting Code TD, the potential for non-linear models to capture intricate relationships cannot be discounted. This is particularly relevant in light of the possibility of incorporating further data in the future and additional model predictors, for instance. Looking at the rest of the ML models and excluding the special case of SVM, non-linear models present more accurate results than linear models in all the calculated metrics.
| Approach | MAPE (%) | MAE | RMSE |
| SARIMAX | 1.00 | 1,539.92 | 2,721.95 |
| SARIMA + LM | 17.67 | 12,921.79 | 14,412.90 |
| ARIMAX | 0.97 | 1,525.96 | 2,646.20 |
| ARIMA + LM | 17.35 | 15,525.21 | 17,312.14 |
| MLR(L) | 4.85 | 4,895.50 | 6,759.43 |
| SVM(NL) | 23.87 | 38,929.00 | 39,904.36 |
| XGB(NL) | 1.72 | 3,006.66 | 4,686.55 |
| RF(NL) | 1.44 | 2,242.57 | 3,878.84 |
| SGD(L) | 5.04 | 5,494.12 | 7,275.05 |
| L1(L) | 5.18 | 5,431.16 | 7,112.04 |
| L2(L) | 4.86 | 4,898.60 | 6,709.81 |
| Approach | MAPE (%) | MAE | RMSE |
| SARIMAX | 1.41 | 2,078.64 | 3,145.27 |
| SARIMA + LM | 13.32 | 13,871.48 | 15,842.22 |
| ARIMAX | 1.65 | 2,349.81 | 3,469.03 |
| ARIMA + LM | 15.02 | 17,586.97 | 19,400.41 |
| MLR(L) | 4.62 | 3,978.65 | 5,567.30 |
| SVM(NL) | 23.12 | 27,822.99 | 28,620.02 |
| XGB(NL) | 5.80 | 4,600.92 | 5,950.18 |
| RF(NL) | 1.90 | 2,572.85 | 4,202.80 |
| SGD(L) | 4.89 | 4,478.40 | 5,956.63 |
| L1(L) | 5.09 | 4,609.15 | 6,067.31 |
| L2(L) | 4.50 | 3,706.95 | 5,192.83 |
Following the set of defined research questions, we wanted to analyze the impact of addressing the seasonality effect on the prediction performance of TSA models through our RQ3. Thus, we conducted the prediction analysis described in Section 3.6 with SARIMAX and SARIMA+LM models. The process was similar to RQ1 and used the same data for this RQ. Figure 5 shows the unified graphical depiction of the predictive performance results from the adopted performance metrics. To avoid overwhelming the reader with too many plots, we displayed the results for the seasonally adjusted models along with their former counterpart models ARIMAX and ARIMA+LM. The displayed results demonstrate a systematic difference in prediction performance among the multivariate single model approach, i.e. ARIMAX and SARIMAX, compared to that of the multivariate combined model approach, i.e. ARIMA+LM and SARIMA+LM. Therefore, we can affirm that the superiority in prediction performance demonstrated on ARIMAX compared to ARIMA+LM is maintained with SARIMAX and SARIMA+LM when adjusting the models for capturing seasonality. Figure 5 demonstrates a clear superiority of SARIMAX over SARIMA+LM given biweekly as well as monthly serialized time series data. Hence, we can affirm that given the adjustment of the existing TSA models for Code TD prediction, the SARIMAX model overcomes the predictive performance of the SARIMA+LM model. Precisely, considering for instance the results for MAE and RMSE the difference regarding the prediction error between the SARIMAX and SARIMA+LM stand in 11,382 and 11,691 within the biweekly data, and 11,792 and 12,697 within the monthly data respectively, thus denoting a remarkable difference among both approaches. Answering to RQ3, the displayed results provide slightly similar results between SARIMAX and ARIMAX models. Results for biweekly data provided better prediction performance for ARIMAX compared to SARIMAX while the opposite happened for monthly data. MAPE results provided worse prediction performance for SARIMA+LM than for ARIMA+LM when using biweekly data. However, SARIMA+LM overtook ARIMA+LM for the MAPE results when using monthly data, as well as in the other prediction performance evaluation metric results. Therefore, we can affirm that adjusting the TSA models to capture the seasonality effects of the data improves their prediction performance to some extent.
Finally, we compared the previously confronted SARIMAX and SARIMA+LM models against the adopted ML algorithms commonly used in the field. Tables 4 and 5 present the prediction performance results for SARIMAX and SARIMA+LM models compared to the adopted ML models. As expected from the results obtained for RQ3, where SARIMAX provided similar prediction results as ARIMAX when using biweekly data and overtook ARIMAX when using monthly data, in this case, the SARIMAX model also demonstrated a clear superiority among the rest of the compared models, thus providing a more accurate prediction outcome for future Code TD observations. It is worth noticing that capturing the seasonality effect did not change the key results observed for RQ2. In fact, SARIMAX did not improve the results of ARIMAX model using biweekly data, and the improvement experimented compared to ARIMAX with monthly resulted to be low. Moreover, RF demonstrated to be the second-best prediction model. Therefore, even if the results suggest the suitability for linear models, non-linear models may also offer the potential for capturing more complex relationships. Looking at the rest of the ML models and excluding the special case of the SVM model, non-linear models present more accurate results than linear models in all the calculated metrics.
5 Discussion
The objective of the presented study was to analyse the impact of temporal factors in Code TD prediction. Code TD was approximated by leveraging the SQALE index metric calculated by SQ, and the predictions were made including the set of different types of code smell issues in the projects within the prediction models. The adoption of TSA prediction models enabled the empirical demonstration of the suitability of time-dependent techniques for Code TD prediction, as well as a comparative analysis with commonly used ML prediction algorithms. Consequently, 11 distinct prediction models were constructed for each project. The analysis included a comprehensive data collection, preprocessing, and examination of 31 open-source Java projects. However, in order to ensure the integrity and fairness of the results, only the outcomes for the 14 projects that provided complete results with the adopted prediction models were considered.
The study was divided into four research questions. The first research question compared the performance of two ARIMA-founded time-dependent approaches already implemented in the academic literature. Following our initial assumption, the ARIMAX model previously proposed by by Mathioudaki et al. [11] outperformed the combination model ARIMA+LM proposed by Zozas et al. [12]. Thus, our initial takeaway depicted the superiority of ARIMAX over ARIMA+LM model when predicting Code TD. These results suggest that, given a time-dependent prediction context, multivariate models which include and simultaneously predict the independent variables as well within the prediction model such as ARIMAX succeed on providing robust results as shown in Figure 5.
Furthermore, in the second research question we proved the robustness reported with ARIMAX by comparing its Code TD prediction performance against a set of the most common ML algorithms adopted in the SE field. The results provided in Tables 5 and 4 corroborated this evidence. These results suggest the necessity of adopting models that consider the time dependence factor in the dependent variable to be predicted. However, it is worth noticing the close results reported by models such as Random Forest or Extreme Gradient Boost, prediction algorithms widely used in academia as well as in industry. Given the non-linear distributional assumption of these models, and the oppositely linear distributional assumption, which is the foundation for TSA models such as ARIMAX, leveraging the combination of these models through ensemble learning approaches might further improve the prediction performance reported in this study.
Moreover, we analyzed the impact of addressing the seasonal component within the time series data on the obtained results. For that, we leveraged the model extensions provided in ARIMA-founded models to analyze the results for SARIMAX and SARIM+LM models. As the takeaway of the performed analysis, we observed slight improvements in the initially adopted TSA models, thus depicting a small impact on the prediction performance from controlling the seasonality component. This evidence suggests a small dependence on seasonality from the perspective of the OS software development community, which might not follow structured development schedules as compared to the software development activity performed in the industry.
As the key takeaway of this study, we can confirm that time-dependent models are applicable and competitive in TD prediction, not only at a single project level but also at an aggregated level as we have demonstrated. Further research is still needed to understand the methodology required to prepare the most suitable data for TSA models. It has been demonstrated that if the data fits well in the TSA model, this latter one can provide high prediction performance. Therefore, this study suggests the adoption of periodic software quality analysis practices within the project development process, thus providing a thorough monitorization of metrics such Code TD and better quality data for prediction performance improvement. Similarly, it is important to note at this stage that when we make our predictions, these values are the obtained most likely estimates, therefore we are not comparing our real values against the result of a simulation or an execution.
TSA models can have a considerable prediction performance, but there is a need for further research in modeling TD prediction while considering time as a relevant factor.
6 Threats to Validity
Construct validity
Some projects have been affected by only a few code smell rule issues, so we decided to exclude them as their independent variables were considered uninformative. Additionally, the irregularity of the commit data can cause missing data for the generated periodical time series. The threat was minimized by linearly interpolating values for the missing periods. Approximations cannot be as accurate as using real data, therefore we could expect different results if periodically complete real data existed.
Internal validity
We selected one of the metrics that depict technical debt, i.e. SQALE index as the variable of interest, and the set of code smell types that SQ reports as the reason for the given SQALE index. However, we did not consider collecting additional independent variables that might affect the results. Therefore, we could expect different results if another choice of independent variables had been made.
External validity
The study subjects were mature open source projects written in Java and fulfilling the rigorous quality standards of ASF. The included projects represent a wide spectrum of application domains, including external libraries, frameworks, web utilities, and substantial computational infrastructures. Hence, the obtained results can be generalized to projects with similar characteristics, and therefore non-mature projects, retired projects, and projects not using SonarQube are excluded.
Conclusion validity
This study applied commonly known statistical and ML techniques. During the data analysis, we ensured the assumptions of the techniques were fulfilled. However, a low number of data points in some projects may reduce the prediction power of the used models. Additionally, the ARIMA + LM model may not capture the time-dependent nature of the SQALE index, as its prediction is performed through a Multiple Linear Regression model, which does not consider the time dependence factor. Regarding the reliability of the measures, the measurements of the variables were collected from the dataset, and the construction of the final set of model variables was made through an automated process. For the sake of reliability on the results, we excluded the collected projects which did not allow the algorithms from the adopted models to converge within the model fitting. Therefore, we report only the aggregated results from the projects which provided results for the entire set of considered models in this study.
7 Related Work
In this section, we report the related work including the current status of Technical Debt (TD) Prediction.
Tsoukalas et al. [56] investigated the effectiveness of ML techniques in modeling and predicting the evolution of TD. Specifically, they carried out an extensive investigation using a dataset they compiled, consisting of weekly snapshots of fifteen open-source software projects spanning three years. This dataset is augmented with software-related metrics extracted from two commonly used static analysis tools, which serve as predictors for TD. Using these identified predictors, they develop a range of TD predicting models employing well-known ML algorithms. These models were then tested against various timeframes for prediction. The findings from their analysis revealed that linear regularization models aptly capture and yield meaningful predictions for TD evolution over short prediction periods. On the other hand, non-linear random forest regression outperforms linear models over longer prediction horizons. In a second work, Tsoukalas et al. [9] adopted ML models to classify TD classes (High/Not-High TD), analyzing 25 open source projects. The achieved results had an accuracy estimation of approximately 0.79, with a recall of 0.86.
Mathioudaki et al. [44] explored the application of Deep Learning to enhance the precision of long-term TD prediction. To achieve this, they created, assessed, and juxtaposed Deep Learning models employing a dataset encompassing five prominent real-world software applications sourced from the Technical Debt dataset version 1 [24]. Results revealed that Deep Learning techniques yield TD prediction models that display commendable predictive accuracy, extending up to 150 steps ahead in future prediction.
Aversano et al. [45] investigated the potential utility of software system quality metrics in accurately predicting TD. To achieve this, information concerning quality metrics from 8 distinct open-source software systems was examined. These metrics were then fed into various ML algorithms to predict technical debt. The outcomes demonstrated strong predictive performance, and the suggested approach offers a valuable method for comprehending the practical aspects of the technical debt phenomenon.
Focusing on the works that adopted TSA models to predict TD, we concentrated on two papers. Mathioudaki et al. [11] investigated the predictive capabilities of TSA models for TD prediction. They aimed to determine whether incorporating independent variables as predictors, known as TD predictors, into ARIMAX models [23, p. 451] could yield more precise TD predictions compared to conventional Autoregressive Integrated Moving Average (ARIMA) models. Their investigation was conducted using five datasets that captured the historical evolution of software metrics derived from static code analysis across five long-standing projects. They generated predictions using both ARIMA and ARIMAX models for various time horizons on this data. Results yielded a clear conclusion. Across the examined open-source software projects, the accuracy of ARIMAX models exceeded that of ARIMA models by a significant margin.
More recently, Zozas et al. [12] combined two models for predicting Technical Debt Principal: Supervised Backwards Stepwise Regression, and ARIMA. Based on the application of Backwards Stepwise Regression they obtained the most significant predictors to describe the TD response variable. They performed univariate TSA prediction through ARIMA for each of the selected predictors to predict future values, and then they used the previously developed regression model to estimate the future values of the response variable.
Comparing our work with Mathioudaki et al. [11], they considered the same data set we did, but they only included 5 projects out of 33, while we investigated the entire dataset. Moreover, in our work, instead of conducting a comparison between univariate and multivariate TSA, we aim to directly perform a comparison between different multivariate TSA approaches to investigate which approach obtains better prediction performance. The selected predictors will be used to predict the value of the SQALE index, which will be our target value. Our approach relies on performing different multivariate applications of the ARIMA model, and for testing its performance we aim to split the data for training and testing the model through the Walk-Forward Train-Test approach.
Furthermore, as we believe that ML models should not be trained with just one feature but the exact amount for the sake of the quality of that model, we compare in this case the performance of the mentioned multivariate TSA models against a set of chosen ML models based on previous research.
8 Conclusions
We conducted an empirical study to compare two TSA approaches, their respecting version adjusted to treat the seasonality component and seven ML algorithms for Code TD prediction as SQALE index computed by SQ.
The compared methods were trained using Code TD observation data that was serialized into biweekly and monthly periodicity levels from 14 open-source Java projects. The comparison denoted a clear superiority of the ARIMAX model compared to the rest of the adopted models when training the models with biweekly time series, while SARIMAX provided better results with monthly time series. However, results provided by SARIMAX suggested that the impact from seasonality might not improve substantially the predictive performance of ARIMAX, thus demonstrating that the main improvement between ML algorithms’ performance and TSA models is generated by the ARIMAX model. Therefore, our findings indicate that, while the seasonality factor might not strongly impact the predictive performance of time-dependent prediction techniques, considering the time-dependence factor improves the prediction results compared to that of time-agnostic techniques. This and the required data serialization performed in this study in order to analyse time series data suggest the potential benefit performing sequential software quality analyses in software projects can have in order to improve the Code TD prediction performance.
In future work, we aim to analyze the implementation of transformer-driven prediction techniques. The adaptation of TSA techniques with these models provides great modelling ability for long-range dependencies and interactions in sequential data and TSA models. This will further contribute to the research community with advanced techniques for predicting and therefore anticipating Code TD.
Appendix: Parameter tuning of the Box-Jenkings models
In this appendix, we describe the mathematical definitions of the model parameters composing the TSA models included in the work. The adopted models stand as modifications of the Box-Jenking model, better known as ARIMA [34]. This model is a univariate TSA model composed of the autoregressive component, moving average component, and differencing component, as explained below.
8.1 Autoregression (AR) order (p / P parameter)
Seasonality adjustment occurs when the autoregressive pattern occurs at multiple lags of the defined periodicity (m). Thus a seasonal autoregression (SAR) process is defined by
where is defined as .
8.2 Moving Average (MA) order (q / Q parameter)
where and is defined as . Every MA(q) process is stationary, but there can be two different MA(q) processes with the same autocovariance [57]. We avoid this issue by enforcing invertible MA processes (see details in Section 3.7).
As before, a seasonality adjustment is performed when the moving average pattern occurs at multiple lags of the defined periodicity (m). Thus seasonal moving average (SMA) processes are defined by
where is defined as .
8.3 Differecing order (d / D parameter)
Among the existing approaches to reach stationarity, models such as ARIMA apply differencing repeatedly to the series of data until the differenced observations resemble a realization of stationary time series, required for making accurate forecasting [57]. Considering for instance an original time series we can define the first differencing order (d = 1) as:
Hence, the differencing operator or lag-d can be formulated as

### v2

@@ Appendix B Replication of the study
results from the ARIMAX model with biweekly time series data
| Biweekly periods | Mean | Median | Max | Min | Variance |
| 1 | 1.44 | 0.12 | 17.82 | 0 | 13.06 |
| 2 | 1.83 | 0.28 | 26.56 | 0 | 25.25 |
| 3 | 2.14 | 0.31 | 28.83 | 0 | 29.38 |
| 4 | 2.96 | 0.4 | 47.36 | 0 | 73.58 |
| 5 | 3.51 | 0.47 | 58.31 | 0 | 110.16 |
| 6 | 4.41 | 0.7 | 65.47 | 0 | 142.34 |
| 7 | 5.09 | 0.87 | 70.46 | 0 | 171.62 |
| 8 | 5.62 | 0.83 | 74.1 | 0 | 202.68 |
| 9 | 6.09 | 0.96 | 76.85 | 0 | 224.2 |
| 10 | 5.46 | 0.92 | 78.65 | 0 | 216.74 |
| 11 | 3.0 | 0.73 | 18.54 | 0 | 21.95 |
| 12 | 3.15 | 0.73 | 19.81 | 0 | 24.43 |
| 13 | 3.29 | 0.76 | 20.85 | 0 | 26.79 |
| 14 | 3.43 | 0.82 | 21.72 | 0 | 28.94 |
| 15 | 3.54 | 0.97 | 22.46 | 0 | 30.69 |
| 16 | 3.69 | 1.1 | 23.11 | 0 | 32.25 |
| 17 | 3.84 | 1.22 | 23.68 | 0 | 33.94 |
| 18 | 4.06 | 1.34 | 24.19 | 0 | 35.83 |
| 19 | 4.01 | 1.39 | 24.76 | 0 | 37.72 |
| 20 | 3.84 | 1.3 | 25.28 | 0 | 38.52 |
| 21 | 4.03 | 1.46 | 25.75 | 0.01 | 40.71 |
| 22 | 4.21 | 1.5 | 26.18 | 0.02 | 43.03 |
| 23 | 4.39 | 1.47 | 26.58 | 0.04 | 45.21 |
| 24 | 4.56 | 1.44 | 26.95 | 0.06 | 47.36 |
| 25 | 4.72 | 1.39 | 27.29 | 0.06 | 49.5 |
| 26 | 4.88 | 1.35 | 27.61 | 0.06 | 51.76 |
| 27 | 5.04 | 1.49 | 27.9 | 0.06 | 53.88 |
| 28 | 4.78 | 1.48 | 28.16 | 0.06 | 54.26 |
| 29 | 4.88 | 1.56 | 28.41 | 0.06 | 55.73 |
| 30 | 4.97 | 1.6 | 28.63 | 0.06 | 56.95 |
| 31 | 5.06 | 1.57 | 28.84 | 0.07 | 58.04 |
| 32 | 4.83 | 1.17 | 29.02 | 0.07 | 59.27 |
| 33 | 4.89 | 1.15 | 29.2 | 0.07 | 59.88 |
| 34 | 4.94 | 1.12 | 29.35 | 0.07 | 60.4 |
| 35 | 4.98 | 1.16 | 29.5 | 0.08 | 60.62 |
| 36 | 5.03 | 1.2 | 29.63 | 0.09 | 60.74 |
| 37 | 5.08 | 1.25 | 29.8 | 0.1 | 61.25 |
| 38 | 5.33 | 1.54 | 29.99 | 0.11 | 63.82 |
| 39 | 5.39 | 1.6 | 30.21 | 0.13 | 64.53 |
| 40 | 5.46 | 1.67 | 30.45 | 0.13 | 65.33 |
| 41 | 5.53 | 1.77 | 30.72 | 0.13 | 66.23 |
| 42 | 5.6 | 1.86 | 31.0 | 0.14 | 67.19 |
| 43 | 5.67 | 1.94 | 31.21 | 0.14 | 67.99 |
| 44 | 5.74 | 2.03 | 31.39 | 0.14 | 68.76 |
| 45 | 5.81 | 2.12 | 31.56 | 0.15 | 69.55 |
| 46 | 5.87 | 2.21 | 31.56 | 0.15 | 69.85 |
| 47 | 5.93 | 2.28 | 31.55 | 0.15 | 70.15 |
| 48 | 5.99 | 2.27 | 31.54 | 0.16 | 70.45 |
| 49 | 6.06 | 2.28 | 31.54 | 0.16 | 70.79 |
| 50 | 4.91 | 1.61 | 23.98 | 0.17 | 40.79 |
| 51 | 4.97 | 1.64 | 24.08 | 0.17 | 41.23 |
| 52 | 5.04 | 1.68 | 24.19 | 0.17 | 41.69 |
| 53 | 5.1 | 1.71 | 24.29 | 0.18 | 42.09 |
| 54 | 5.17 | 1.75 | 24.39 | 0.18 | 42.49 |
| 55 | 5.24 | 1.79 | 24.48 | 0.18 | 42.87 |
| 56 | 5.32 | 1.82 | 24.57 | 0.19 | 43.23 |
| 57 | 5.39 | 1.86 | 24.66 | 0.19 | 43.62 |
| 58 | 5.22 | 1.82 | 24.74 | 0.19 | 44.95 |
| 59 | 5.5 | 1.93 | 24.81 | 0.2 | 46.92 |
| 60 | 5.57 | 1.97 | 24.89 | 0.2 | 47.28 |
| 61 | 5.64 | 2.01 | 24.96 | 0.21 | 47.62 |
| 62 | 5.91 | 2.9 | 25.04 | 0.21 | 49.97 |
| 63 | 5.5 | 2.13 | 25.11 | 0.21 | 49.02 |
| 64 | 5.57 | 2.21 | 25.18 | 0.22 | 49.36 |
| 65 | 5.64 | 2.29 | 25.25 | 0.22 | 49.76 |
| 66 | 5.74 | 2.29 | 25.32 | 0.22 | 53.49 |
| 67 | 5.8 | 2.36 | 25.38 | 0.23 | 53.94 |
| 68 | 5.87 | 2.42 | 25.45 | 0.23 | 54.41 |
| 69 | 5.93 | 2.48 | 25.51 | 0.24 | 54.83 |
| 70 | 5.99 | 2.55 | 25.57 | 0.24 | 55.35 |
| 71 | 5.91 | 2.52 | 25.64 | 0.24 | 59.57 |
| 72 | 5.97 | 2.58 | 25.7 | 0.25 | 60.07 |
results from the SARIMAX model with monthly time series data
| Monthly periods | Mean | Median | Max | Min | Variance |
| 1 | 0.3 | 0.08 | 1.52 | 0 | 0.18 |
| 2 | 0.34 | 0.14 | 1.2 | 0 | 0.15 |
| 3 | 0.5 | 0.25 | 2.1 | 0 | 0.32 |
| 4 | 0.61 | 0.31 | 2.72 | 0 | 0.51 |
| 5 | 0.73 | 0.39 | 3.32 | 0 | 0.76 |
| 6 | 1.08 | 0.52 | 6.51 | 0 | 2.42 |
| 7 | 1.38 | 0.57 | 10.37 | 0 | 5.36 |
| 8 | 1.61 | 0.64 | 13.2 | 0 | 8.43 |
| 9 | 1.79 | 0.73 | 15.39 | 0 | 11.4 |
| 10 | 1.98 | 0.73 | 17.43 | 0 | 14.37 |
| 11 | 2.15 | 0.72 | 19.07 | 0 | 17.04 |
| 12 | 2.33 | 0.72 | 20.4 | 0 | 19.46 |
| 13 | 2.49 | 0.72 | 21.5 | 0 | 21.67 |
| 14 | 2.63 | 0.72 | 22.41 | 0 | 23.69 |
| 15 | 2.76 | 0.72 | 23.2 | 0 | 25.6 |
| 16 | 2.89 | 0.72 | 23.89 | 0 | 27.21 |
| 17 | 3.03 | 0.71 | 24.49 | 0 | 28.83 |
| 18 | 3.15 | 0.71 | 25.03 | 0 | 30.44 |
| 19 | 3.29 | 0.71 | 25.61 | 0 | 32.2 |
| 20 | 3.44 | 0.78 | 26.13 | 0 | 33.6 |
| 21 | 3.59 | 0.76 | 26.59 | 0.01 | 34.88 |
| 22 | 3.73 | 0.83 | 27.01 | 0.02 | 36.16 |
| 23 | 3.88 | 0.89 | 27.4 | 0.03 | 37.61 |
| 24 | 4.03 | 0.96 | 27.74 | 0.05 | 39.06 |
| 25 | 4.17 | 1.03 | 28.05 | 0.08 | 40.56 |
| 26 | 4.31 | 1.16 | 28.34 | 0.1 | 42.21 |
| 27 | 4.45 | 1.41 | 28.6 | 0.13 | 43.8 |
| 28 | 4.11 | 1.34 | 28.85 | 0.16 | 42.31 |
| 29 | 4.19 | 1.36 | 29.07 | 0.19 | 43.19 |
| 30 | 4.26 | 1.39 | 29.28 | 0.23 | 43.87 |
| 31 | 4.31 | 1.42 | 29.46 | 0.27 | 44.55 |
| 32 | 3.96 | 1.33 | 29.63 | 0.3 | 43.58 |
| 33 | 4.01 | 1.31 | 29.79 | 0.31 | 44.0 |
| 34 | 4.06 | 1.29 | 29.93 | 0.32 | 44.36 |
| 35 | 4.11 | 1.27 | 30.06 | 0.33 | 44.77 |
| 36 | 4.19 | 1.28 | 30.19 | 0.34 | 45.35 |